### v1

# Extremal simplicial distributions on cycle scenarios with arbitrary outcomes

Abstract
Cycle scenarios are a significant class of contextuality scenarios, with the Clauser-Horne-Shimony-Holt (CHSH) scenario being a notable example. While binary outcome measurements in these scenarios are well understood, the generalization to arbitrary outcomes remains less explored, except in specific cases. In this work, we employ homotopical methods in the framework of simplicial distributions to characterize all contextual vertices of the non-signaling polytope corresponding to cycle scenarios with arbitrary outcomes. Additionally, our techniques utilize the bundle perspective on contextuality and the decomposition of measurement spaces. This enables us to extend beyond scenarios formed by gluing cycle scenarios and describe contextual extremal simplicial distributions in these generalized contexts.
1 Introduction
Simplicial distributions [1] are a new framework for studying contextuality that extends the sheaf-theoretic framework of Abramsky–Brandenburger [2] and the cohomological framework of Okay et al. [3]. Measurements and outcomes are represented by combinatorial models of spaces known as simplicial sets, and the non-signaling conditions are encoded topologically. The theory of simplicial distributions is an extension of the theory of non-signaling distributions, which utilizes topological ideas and tools for studying the polytopes of non-signaling distributions and the Bell polytopes describing the classical region [4]. In this paper, we use homotopical techniques to identify the extremal distributions on cycle scenarios with arbitrary outcomes that extend earlier results in the literature. Our techniques generalize to more sophisticated measurement spaces obtained by gluing cycle scenarios.
A simplicial distribution consists of a family of distributions $p_{\sigma}$ parametrized by the simplices of the measurement space. Here $\sigma$ is an $n$-dimensional measurement (or context), and the distribution $p_{\sigma}$ is on the set of $n$-dimensional outcomes. These distributions are related by the topological non-signaling conditions imposed by the simplicial structure of the measurement and outcome spaces. In this paper, we consider a measurement space that is $1$-dimensional, e.g., a directed graph, whose vertices represent measurements and edges consist of pairs of measurements that can be simultaneously performed. A convenient way to represent these simplicial distribution is as a family of $d\times d$ matrices with entries $p_{\sigma}^{ab}$ for each edge $\sigma$. The sum of the rows of this matrix gives the marginal distribution at the source vertex and the sum of the columns at the target vertex. A cycle scenario has the measurement space $C^{(n)}$ consisting of $n$-edges glued to form a disk’s boundary. For example, $n=4$ corresponds to the famous Clauser–Horne–Shimony–Holt (CHSH) scenario [5]. Our main result is a characterization of the contextual vertices of cycle scenarios with arbitrary outcomes in $\mathbb{Z}_{d}=\{0,1,\cdots,d-1\}$. A $k$-order cycle distribution, where $1\leq k\leq d$, is a particular type of distribution whose non-zero probabilities are of the form $1/k$ (Definition 4.1).
Theorem.
A simplicial distribution on the cycle scenario $C^{(n)}$ is a contextual vertex if and only if it is a $k$-order cycle distribution for some $k\geq 2$.
Our main method of approach follows the homotopical methods introduced in [6]. There are two ways to extend these results:
  • The bundle approach of [7] allows us to construct scenarios where measurements can assume outcomes in different sets. Corollary 4.6 extends our theorem above to the case of arbitrary outcomes not necessarily uniform over each measurement, extending the results of [8].
  • We analyze how gluing two measurement spaces affects the vertices of the associated polytope of simplicial distributions. Theorem 5.3 allows us to extend our results to scenarios obtained by gluing cycle scenarios.
The structure of the paper is as follows. In Section 2, we introduce background material for simplicial sets and homotopical tools for analyzing faces of distribution polytopes. Section 3 is about bundle scenarios and their application to extend our analysis to larger sets of outcomes, as demonstrated in Proposition 3.5. In Section 4, we prove our main theorem highlighted above. Gluings of cycle scenarios are analyzed in Section 5 along with interesting examples.
2 Homotopical vertices of faces
We begin by recalling basic definitions from the theory of simplicial distributions [1]. Then in Section 2.2 we introduce a new notion called the vertex support based on a preorder between simplicial distributions. We use this notion to provide a characterization of vertices. In Sections 2.3 and 2.4 we recall results from [6] on the homotopical characterization of strong contextuality.
2.1 Simplicial distributions
The theory of simplicial distributions [1] is a combinatorial framework for describing distributions on a simplicial scenario $(X,Y)$ consisting of
  • a measurement space $X$, and
  • an outcome space $Y$.
A space in this framework is represented by a simplicial set $X$ consisting of the data of a set of simplicies $X_{0},X_{1},\cdots,X_{n},\cdots$ and the simplicial structure maps. The structure maps, given by the face maps $d_{i}^{X}:X_{n}\to X_{n-1}$ and degeneracy maps $s_{j}^{X}:X_{n}\to X_{n+1}$, encode how to glue and collapse the simplices, respectively (see, e.g.,[9]). We usually denote the simplicial structure maps by $d_{i}$ and $s_{j}$ omitting the simplicial set from the notation. A simplex is called degenerate if it lies in the image of a degeneracy map. Otherwise, it is called non-degenerate. A non-degenerate simplex is called a generating simplex if it is not a face of another simplex. A map $f:X\to Y$ between two simplicial sets is given by a collection of functions $\{f_{n}:X_{n}\to Y_{n}\}_{n\geq 0}$ compatible with the simplicial structure maps. We will employ the notation $f_{x}=f_{n}(x)$ for $x\in X_{n}$. The set of simplicial set maps will be denoted by $s\mathbf{Set}(X,Y)$.
For a set $U$, let $D(U)$ denote the set of probability distributions, i.e., functions $p:U\to\mathbb{R}_{\geq 0}$ with finite support satisfying $\sum_{u\in U}p(u)=1$. A simplicial distribution on the scenario $(X,Y)$ is a simplicial set map
$p:X\to D(Y).$
Here $D(Y)$ is the simplicial set whose $n$-simplices are given by the set $D(Y_{n})$ of distributions on $Y_{n}$ and the simplicial structure maps are given by marginalization along the structure maps of $Y$:
${d_{i}^{D(Y)}=D(d_{i}^{Y})\;\;\text{ and }\;\;s_{j}^{D(Y)}=D(s_{j}^{Y}).}$
More explicitly, a simplicial distribution consists of a family of distributions
$\{p_{x}\in D(Y_{n}):x\in X_{n}\}_{n\geq 0}$
compatible under the topological non-signaling conditions given by $d_{i}^{D(Y)}p_{x}=p_{d_{i}x}$ and $s_{j}^{D(Y)}p_{x}=p_{s_{j}x}$. We write $p_{x}^{y}$ for the probability $p_{x}(y)$ of observing the outcome $y$ for the measurement $x$. For a simplicial set map $\varphi:X\to Y$ we define the associated deterministic distribution $\delta^{\varphi}:X\to D(Y)$ by $\delta_{Y}\circ\varphi$ where $\delta_{Y}:Y\to D(Y)$ is the canonical map that sends a simplex to the delta distribution peaked at that simplex. The set ${s\mathbf{Set}(X,D(Y))}$ of simplicial distributions on $(X,Y)$ is denoted by $\operatorname{sDist}(X,Y)$.
Typically in applications we impose the following restrictions:
  • $X$ is a finitely generated simplicial set, and
  • $Y$ has finitely many simplices in each dimension.
The first condition means that $X$ has finitely many generating simplices. Under these conditions $\operatorname{sDist}(X,Y)$ is a polytope.
We are primarily interested in the following spaces, which will serve as measurement and outcome spaces, respectively:
  • An $n$-circle is a simplicial set $C^{(n)}$ specified by a sequence of pairwise distinct $1$-simplices $\sigma_{1},\cdots,\sigma_{n}$ satisfying $d_{0}(\sigma_{i})=d_{1}(\sigma_{i+1})$ for every $11\leq i\leq n-$ 1 and $d_{0}(\sigma_{n})=d_{1}(\sigma_{1})$. We will write $v_{i}=d_{1}(\sigma_{i})$.
  • For a set $U$ let $\Delta_{U}$ denote the simplicial set whose $n$-simplices are given by the set $U^{n+1}$ and the simplicial structure maps are given by
    $d_{i}(x_{0},x_{1},\cdots,x_{n})$ $=(x_{0},x_{1},\cdots,x_{i-1},x_{i+1},\cdots,x_{n})$
    $s_{j}(x_{0},x_{1},\cdots,x_{n})$ $=(x_{0},x_{1},\cdots,x_{j-1},x_{j},x_{j},x_{j+1},\cdots,x_{n}).$
Definition 2.1.
An $n$-cycle scenario consists of
  • the measurement space given by an $n$-circle $C^{(n)}$, and
  • the outcomes space $\Delta_{\mathbb{Z}_{d}}$ where $\mathbb{Z}_{d}=\{0,1,\cdots,d-1\}$.
In Figure 1 the $4$-circle measurement space is depicted. When the outcomes are in $\mathbb{Z}_{2}$ the resulting scenario is the famous CHSH scenario.
Figure 1: The $4$-circle measurement space.
More generally, we will consider $1$-dimensional measurement spaces, i.e., basically directed graphs. In this case the simplicial set representing the measurement space is specified by its 00-simplices and $1$-simplices together with the face maps between them. We will write a distribution on an edge $\sigma$ as a $d\times d$-matrix $M$ where $M_{ab}=p^{ab}$. For convenience we index the rows and columns of such a matrix by $10,\cdots,d-$ 1. The marginals on the initial vertex and the terminal vertex are given by the sum of the rows and columns, respectively.
2.2 The vertex support
Next, we introduce a preorder on the set of simplicial distributions and an associated notion called the vertex support. This is used to provide a new characterization of vertices.
Definition 2.2.
The support of a simplicial distribution $p:X\to D(Y)$ is defined by
$\operatorname{supp}(p)=\{\varphi\in s\mathbf{Set}(X,Y):\,p_{x}^{\varphi_{x}}\neq 0,\;\forall x\in X_{n},\,n\geq 0\}.$
We say $p:X\to D(Y)$ is strongly contextual if $\operatorname{supp}(p)$ is empty.
Note that to have $\varphi\in\operatorname{supp}(p)$ it suffices that $p_{x}^{\varphi_{x}}\neq 0$ for every generating simplex $x$ of $X$.
Definition 2.3.
Given two simplicial distributions $p,q:X\to D(Y)$ we write $q\preceq p$ if $q_{x}^{y}\neq 0$ implies $p_{x}^{y}\neq 0$ for every $n\geq 0$, $x\in X_{n}$, and $y\in Y_{n}$.
The relation $\preceq$ is a preorder on $\operatorname{sDist}(X,Y)$. It suffices to verify the condition for this relation only on the generating simplices of $X$. Recall that a simplicial distribution $p$ is called extremal if $p=\alpha q+(1-\alpha)\tilde{q}$ and $0<\alpha\leq 1$ implies $p=q$. Extremal simplicial distributions are also called vertices.
Definition 2.4.
Given a simplicial distribution $p:X\to D(Y)$. We define the vertex support $\operatorname{Vsupp}(p)$ to be the set of vertices $q\in\operatorname{sDist}(X,Y)$ satisfying $q\preceq p$.
The notion of vertex support is a natural extension of the notion of support. Note that $\varphi\in\operatorname{supp}(p)$ if and only if $\delta^{\varphi}\preceq p$. Since every deterministic distribution is a vertex we have $\{\delta^{\varphi}~{}|~{}\varphi\in supp(p)\}\subseteq\operatorname{Vsupp}(p)$.
Lemma 2.5.
Given simplicial distributions $p,q:X\to D(Y)$, we have $q\preceq p$ if and only if there exists a simplicial distribution $\tilde{p}:X\to D(Y)$ and $\alpha\in(0,1]$ such that $p=\alpha q+(1-\alpha)\tilde{p}$.
Proof.
Assume $q\preceq p$. We define $\alpha$ to be the minimum of the following set of real numbers
$\bigcup_{n\geq 0}\{\frac{p_{x}^{y}}{q_{x}^{y}}~{}|~{}x\in X_{n}~{}\text{a generator},y\in Y_{n},~{}\text{and}~{}q_{x}^{y}\neq 0\}.$
Firstly, this minimum exists because the set is finite and $\alpha>0$ because $q\preceq p$. On the other hand, the condition that $\sum_{y\in Y_{n}}p_{x}^{y}=1=\sum_{y\in Y_{n}}q_{x}^{y}$ for every $x\in X_{n}$, implies that for every $x\in X_{n}$ there is $y\in Y_{n}$ such that $q_{x}^{y}\neq 0$ and $p_{x}^{y}\leq q_{x}^{y}$, which means that $\alpha\leq 1$. If $\alpha=1$ we conclude that $p=q$, in which case $p=1\cdot p+(1-1)p$. For the rest we assume that $\alpha<1$. Then we define the simplicial distribution $\tilde{p}:X\to D(Y)$ by
$\tilde{p}_{x}^{y}=\frac{p_{x}^{y}-\alpha q_{x}^{y}}{1-\alpha}$
for every generator $x\in X_{n}$ and $y\in Y_{n}$. We have
$\sum_{y\in Y_{n}}\tilde{p}_{x}^{y}=\sum_{y\in Y_{n}}\frac{p_{x}^{y}-\alpha q_{x}^{y}}{1-\alpha}=\frac{1}{1-\alpha}(\sum_{y\in Y_{n}}p_{x}^{y}-\alpha\sum_{y\in Y_{n}}q_{x}^{y})=\frac{1}{1-\alpha}(1-\alpha)=1.$
The map $\tilde{p}$ is simplicial since both $p$ and $q$ are simplicial. Finally, one can check that $p=\alpha q+(1-\alpha)\tilde{p}$. The converse direction is clear. ∎
An immediate consequence of this result is a characterization of vertices.
Corollary 2.6.
A simplicial distribution $p:X\to D(Y)$ is a vertex if and only if $q\preceq p$ implies that $q=p$.
Proposition 2.7.
For a simplicial distribution $p:X\to D(Y)$, the following properties hold:
  1. 1.
    If $p$ is a contextual vertex then $p$ is strongly contextual.
  2. 2.
    If $p$ is strongly contextual then it is contextual.
Proof.
To prove (1) suppose that $\varphi\in\operatorname{supp}(p)$, i.e., $\delta^{\varphi}\preceq p$. By Corollary 2.6 this implies that $p=\delta^{\varphi}$, a contradiction since $p$ is contextual. Part (2) is proved in [10, Pro. 2.12]. ∎
2.3 Null-homotopy
Simplicial sets provide a combinatorial framework for homotopy theory [11]. In this section, we will be restricting our attention to the notion of null-homotopy and defining it in a way best suited for our purposes. First, we introduce an important construction.
Definition 2.8.
([12]) The décalage of a simplicial set $Y$ is the simplicial set $\text{Dec}^{0}(Y)$ obtained by shifting the simplices of $Y$ down by one degree, i.e., $\text{Dec}^{0}(Y)_{n}=Y_{n+1}$, and forgetting the first face and degeneracy maps.
We have a simplicial map $d_{0}:\text{Dec}^{0}(Y)\to Y$ that sends a simplex to its 00-th face. For $Y$ we will take the nerve space $N\mathbb{Z}_{d}$. The set of $n$-simplices of the nerve space consists of $n$-tuples of elements $(a_{1},a_{2},\cdots,a_{n})\in\mathbb{Z}_{d}^{n}$ together with the face and the degeneracy maps:
$d_{i}(a_{1},\cdots,a_{n})$ =$=\left\{\begin{array}[]{ll}(a_{2},\cdots,a_{n})&i=0\\ (a_{1},\cdots,a_{i}+a_{i+1},\cdots,a_{n})&0<i<n\\ (a_{1},\cdots,a_{n-1})&i=n\end{array}\right.$
$s_{j}(a_{1},\cdots,a_{n})$ $=(a_{1},\cdots,a_{j},0,a_{j+1},\cdots,a_{n}).$
Definition 2.9.
A simplicial map $\varphi:X\to N\mathbb{Z}_{d}$ is called null-homotopic if there exists a simplicial map $\psi:X\to\text{Dec}^{0}(N\mathbb{Z}_{d})$ such that $d_{0}\circ\psi=\varphi$.
Next we provide an alternative description for the décalage of the nerve space.
Lemma 2.10.
There is an isomorphism of simplicial sets
$\Delta_{\mathbb{Z}_{d}}\xrightarrow{\cong}\text{Dec}^{0}(N\mathbb{Z}_{d})$
defined in degree $m$ by sending $(a_{0},a_{1},\cdots,a_{m})$ to the tuple $(a_{0},a_{1}-a_{0},a_{2}-a_{1},\cdots,a_{m}-a_{m-1})$.
Proof.
See [6, Lemma 3.30]. ∎
Therefore the canonical map $d_{0}:\text{Dec}^{0}(N\mathbb{Z}_{d})\to N\mathbb{Z}_{d}$ can be replaced by the map
$\kappa:\Delta_{\mathbb{Z}_{d}}\to N\mathbb{Z}_{d}$ (1)
defined in degree $m$ by sending $(a_{0},a_{1},\cdots,a_{m})$ to $(a_{1}-a_{0},a_{2}-a_{1},\cdots,a_{m}-a_{m-1})$.
Proposition 2.11.
The simplicial map $\varphi:C^{(n)}\to N\mathbb{Z}_{d}$ is null-homotopic if and only if $\sum_{i=1}^{n}\varphi_{\sigma_{i}}=0$.
Proof.
See [6, Proposition 3.4]. ∎
2.4 Vertices of faces
Next, we define the face associated to a deterministic simplicial distribution and show that homotopy can detect its vertices.
Definition 2.12.
Given $\varphi\in s\mathbf{Set}(X,N\mathbb{Z}_{d})$ we define the face $\operatorname{Face}(\varphi)$ at $\varphi$ to be the preimage of $\delta^{\varphi}$ under the map
$D(\kappa)_{\ast}:s\mathbf{Set}(X,D(\Delta_{\mathbb{Z}_{d}}))\to s\mathbf{Set}(X,D(N\mathbb{Z}_{d}))$
where $\kappa$ is as given in (1).
In other words, a simplicial distribution $p:X\to D(\Delta_{\mathbb{Z}_{d}})$ belongs to $\operatorname{Face}(\varphi)$ if and only if it makes the following diagram commute
${D(\Delta_{\mathbb{Z}_{d}})}$${X}$${D(N\mathbb{Z}_{d})}$$\scriptstyle{D(\kappa)}$$\scriptstyle{p}$$\scriptstyle{\delta^{\varphi}}$
Our main result in this section is the homotopical detection of the vertices of faces.
Proposition 2.13.
Given $\varphi\in s\mathbf{Set}(X,N\mathbb{Z}_{d})$, we have the following:
  1. 1.
    Every vertex in $\operatorname{Face}(\varphi)$ is a vertex in $\operatorname{sDist}(X,\Delta_{\mathbb{Z}_{d}})$.
  2. 2.
    If $\varphi:X\to N\mathbb{Z}_{d}$ is not null-homotopic, then every simplicial distribution in $\operatorname{Face}(\varphi)$ is strongly contextual.
Proof.
Part (1) follows from Lemma 2.10 and by part (1) of Proposition 2.21 in [6]. Part (2) follows from Proposition 2.22 in [6]. ∎
Corollary 2.14.
If $\varphi$ is not null-homotopic and $\operatorname{Face}(\varphi)=\{p\}$ then $p$ is a contextual vertex of $\operatorname{sDist}(X,\Delta_{\mathbb{Z}_{d}})$.
3 Bundle scenarios
Bundle scenarios are first introduced in [7] for developing a resource theory of contextuality in the framework of simplicial distributions. In this section we recall the basics of this theory and illustrate how it can be applied to study cycle scenarios with arbitrary outcomes.
3.1 Simplicial distributions on bundles
A bundle scenario consists of a simplicial set map
$f:E\to X$
where
  • $X$ represents the space of measurements, and
  • $E$ represents the space of events where each fiber $f^{-1}(x)$ over a simplex $x\in X_{n}$ represents the set of outcomes for the measurement.
This formalism is introduced in [7], where the notion is more restrictive. There, only certain kinds of simplicial set maps are termed bundle scenarios. However, the theory also works in this generality of arbitrary simplicial maps. With this generality, other interesting cases include principal bundles as scenarios which yield a twisted theory of simplicial distributions [13]. Every scenario $(X,Y)$ in the ordinary sense can be regarded as a bundle scenario given by the projection map ${\text{pr}_{1}:}X\times Y\to{X}$.
A simplicial distribution on the bundle scenario $f:E\to X$ is a family of distributions
$\{p_{x}\in D(f^{-1}(x)):\,x\in X_{n}\}_{n\geq 0}$
satisfying a compatibility condition induced by the simplicial structure of the measurements and the events. More formally, a simplicial distribution $p$ on the bundle scenario $f$ is given by a commutative diagram
${D(E)}$${X}$${D(X)}$$\scriptstyle{D(f)}$$\scriptstyle{p}$$\scriptstyle{\delta_{X}}$
The set of simplicial distributions on $f$ is denoted by $\operatorname{sDist}(f)$. For every simplicial scenario $(X,Y)$ the set $\operatorname{sDist}(X,Y)$ is in bijective correspondence with $\operatorname{sDist}(\text{pr}_{1})$.
Here is an example of a bundle scenario which will be important for us in the next section.
Example 3.1.
Given a $1$-dimensional simplicial set $X$ and a tuple of natural numbers $\vec{m}=(m_{x})_{x\in X_{0}}$, where $m_{x}\geq 2$, we define the $1$-dimensional simplicial set $E(X,\vec{m})$ by setting
  • $E(X,\vec{m})_{0}=\bigsqcup_{x\in X_{0}}\{x\}\times\mathbb{Z}_{m_{x}}$, and
  • $E(X,\vec{m})_{1}=\bigsqcup_{\sigma\in X_{1}}\{\sigma\}\times\mathbb{Z}_{m_{d_{1}(\sigma)}}\times\mathbb{Z}_{m_{d_{0}(\sigma)}}$ where
$d_{1}(\sigma,(i,j))$ $=(d_{1}(\sigma),i)$
$d_{0}(\sigma,(i,j))$ $=(d_{0}(\sigma),j)$
$s_{0}(x,i)$ $=(s_{0}(x),(i,i)).$
The evident projection map induces a bundle scenario
$f_{m}:E(X,\vec{m})\to X.$
3.2 Morphisms of bundle scenarios
The advantage of introducing bundle scenarios lies in the notion of morphisms between them. Given bundle scenarios $f:E\to X$ and $g:E^{\prime}\to{X^{\prime}}$, a morphism $(\pi,\alpha):f\to g$ between them is given by a commutative diagram
[…]
where the left-hand square is a pull-back.
Here is an example of a map between bundle scenarios where $\pi$ is the identity map:
Example 3.2.
Given injective maps $t_{i}:\mathbb{Z}_{d_{1}}\to\mathbb{Z}_{d_{2}}$, $1\leq i\leq n$, we define a simplicial map $T:C^{(n)}\times\Delta_{\mathbb{Z}_{d_{1}}}\to C^{(n)}\times\Delta_{\mathbb{Z}_{d_{2}}}$ by setting
$T(v_{i},a)=(v_{i},t_{i}(a))$
where $a\in\mathbb{Z}_{d_{1}}$ and $v_{i}$ is a vertex of $C^{(n)}$ (see Definition 2.1). To see that this assignment determines the simplicial map, observe that $T(\sigma_{i},(a,b))$ is specified by $(\sigma_{i},(t_{i}(a),t_{i+1}(b)))$, and for every degenerate simplex $\sigma$ of $C^{(n)}$ the image of $(\sigma,(a_{1},\cdots,a_{m}))$ is determined using the degeneracy maps.
Definition 3.3.
Given a morphism $(\pi,\alpha):f\to g$ between two bundle scenarios and a simplicial distribution $p$ on $f$ the push-forward distribution $q=(\pi,\alpha)_{*}(p)$ on $g$ is defined by
$q_{{x^{\prime}}}^{\theta}=\sum_{\gamma:\alpha(\gamma,{x^{\prime}})=\theta}p_{\pi({{x^{\prime}}})}^{\gamma}$
where $\gamma\in E_{n}$ in the sum also satisfies that $f(\gamma)=\pi({{x^{\prime}}})$.
Note that in the case where $\pi$ is the identity map, we have $(\operatorname{Id},\alpha)_{*}(p)=D(\alpha)\circ p$.
Lemma 3.4.
Consider a morphism of bundle scenarios
[…]
where $i$ is injective. A simplicial distribution $q\in\operatorname{sDist}(g)$ lies in the image of $i_{\ast}:\operatorname{sDist}(f)\to\operatorname{sDist}(g)$ if and only if $q_{x}(e)=0$ for every $x\in X_{n}$ and $e\in E_{n}-\operatorname{Im}(i_{n})$.
Proof.
Suppose we have $p\in\operatorname{sDist}(f)$ such that $i_{\ast}(p)=q$. Then for $x\in X_{n}$ and $e\in E_{n}-\operatorname{Im}(i_{n})$ we have
$q_{x}(e)=(i_{\ast}(p))_{x}(e)=D(i_{n})(p_{x})(e)=\sum_{e^{\prime}:i_{n}(e^{\prime})=e}p_{x}(e^{\prime}).$
So if $e\notin\operatorname{Im}(i_{n})$ then $q_{x}(e)=0$.
For the converse, suppose that $q_{x}(e)=0$ for every $x\in X_{n}$ and $e\in E_{n}-\operatorname{Im}(i_{n})$. For $n\geq 0$, we define $p_{n}:X_{n}\to D(E^{\prime}_{n})$ to be $p_{x}(e^{\prime})=q_{x}(i_{n}(e^{\prime}))$ where $x\in X_{n}$ and $e^{\prime}\in E^{\prime}_{n}$. First we prove that $\{p_{n}\}_{n\geq 0}$ form a simplicial map $p:X\to D(E^{\prime})$. Given $x\in X_{n}$ and $e^{\prime}\in E^{\prime}_{n-1}$, we have
$p_{d_{k}(x)}(e^{\prime})=q_{d_{k}(x)}(i_{n-1}(e^{\prime}))=D(d_{k})(q_{x})(i_{n-1}(e^{\prime}))=\sum_{e:\,d_{k}(e)=i_{n-1}(e^{\prime})}q_{x}(e).$
Since $q_{x}(e)=0$ for $e\notin\operatorname{Im}(i_{n})$ and because $i_{n}$ is injective we have
$\sum_{e:\,d_{k}(e)=i_{n-1}(e^{\prime})}q_{x}(e)$ $=\sum_{e^{\prime\prime}:\,d_{k}(i_{n}(e^{\prime\prime}))=i_{n-1}(e^{\prime})}q_{x}(i_{n}(e^{\prime\prime}))$
$=\sum_{e^{\prime\prime}:\,i_{n-1}(d_{k}(e^{\prime\prime}))=i_{n-1}(e^{\prime})}q_{x}(i_{n}(e^{\prime\prime}))$
$=\sum_{e^{\prime\prime}:\,d_{k}(e^{\prime\prime})=e^{\prime}}q_{x}(i_{n}(e^{\prime\prime})).$
On the other hand, we have
$D(d_{k})(p_{x})(e^{\prime})=\sum_{e^{\prime\prime}:\,d_{k}(e^{\prime\prime})=e^{\prime}}p_{x}(e^{\prime\prime})=\sum_{e^{\prime\prime}:\,d_{k}(e^{\prime\prime})=e^{\prime}}q_{x}(i_{n}(e^{\prime\prime})).$
This shows that $p_{d_{k}(x)}=D(d_{k})(p_{x})$. Analogously, one can prove that $p_{s_{k}(x)}=D(s_{k})(p_{x})$. Now, we prove that $D(i)\circ p=q$. Given $x\in X_{n}$ and $e\in E_{n}$, since $i_{n}$ is injective we obtain
$D(i_{n})(p_{x})(e)=\sum_{e^{\prime}:\,i_{n}(e^{\prime})=e}p_{x}(e^{\prime})=\sum_{e^{\prime}:\,i_{n}(e^{\prime})=e}q_{x}(i_{n}(e^{\prime}))=\begin{cases}q_{x}(e)&\text{if}\;\;e\in\operatorname{Im}{(i_{n})}\\ 0&\text{else}.\end{cases}$
Using this we conclude that
$D(f)\circ p{=D(g\circ i)\circ p}=D(g)\circ D(i)\circ p=D(g)\circ q=\delta_{X}.$
Therefore $p\in\operatorname{sDist}(f)$. ∎
Our main result in this section is the preservation of some of the important features of distributions along the push-forward map.
Proposition 3.5.
Consider a morphism of bundle scenarios
[…]
where $i$ is injective. For $p\in\operatorname{sDist}(f)$, we have the following:
  1. 1.
    $p$ is a deterministic distribution if and only if $i_{\ast}(p)$ is a deterministic distribution.
  2. 2.
    $p$ is a vertex if and only if $i_{\ast}(p)$ is a vertex.
  3. 3.
    $p$ is a contextual vertex if and only if $i_{\ast}(p)$ is a contextual vertex.
Proof.
Part (1): Consider a deterministic distribution $\delta^{\varphi}=\delta_{E^{\prime}}\circ\varphi\in\operatorname{sDist}(f)$, where $\varphi:X\to E^{\prime}$ is a section for $f$. In this case $i\circ\varphi$ is a section for $g$ and $i_{\ast}(\delta^{\varphi})=D(i)\circ\delta_{E^{\prime}}\circ\varphi=\delta_{E}\circ i\circ\varphi=\delta^{i\circ\varphi}$.
On the other hand, if $i_{\ast}(p)=\delta^{\psi}$ for some section $\psi$ for $g$, then for every $x\in X_{n}$ we have
$1=\delta^{\psi_{x}}(\psi_{x})=i_{\ast}(p)_{x}(\psi_{x})=D(i_{n})(p_{x})(\psi_{x})=\sum_{e^{\prime}:\,i_{n}(e^{\prime})=\psi_{x}}p_{x}(e^{\prime}).$
Since the map $i_{n}$ is injective, there is a unique $e^{\prime}\in E^{\prime}_{n}$ with $i_{n}(e^{\prime})=\psi_{x}$ and $p_{x}(e^{\prime})=1$. This means that $p$ is a deterministic distribution.
Part (2): Suppose that $p\in\operatorname{sDist}(f)$ is a vertex. If there exist $q,s\in\operatorname{sDist}(g)$ and $10<\alpha<$ 1 such that $i_{\ast}(p)=\alpha q+(1-\alpha)s$, then by Lemma 3.4 we have
$0=i_{\ast}(p)_{x}(e)=\alpha q_{x}(e)+(1-\alpha)s_{x}(e)$
for every $x\in X_{n}$ and $e\in E_{n}-\operatorname{Im}(i_{n})$. This implies that $q_{x}(e)=s_{x}(e)=0$. Again, by Lemma 3.4 we have $\tilde{q}$ and $\tilde{s}$ in $\operatorname{sDist}(f)$ such that $i_{\ast}(\tilde{q})=q$ and $i_{\ast}(\tilde{s})=s$. Therefore, using the convexity of $i_{\ast}$ we obtain
$i_{\ast}(p)=\alpha i_{\ast}(\tilde{q})+(1-\alpha)i_{\ast}(\tilde{s})=i_{\ast}(\alpha\tilde{q}+(1-\alpha)\tilde{s}).$
Since $i_{\ast}$ is injective, we obtain that $p=\alpha\tilde{q}+(1-\alpha)\tilde{s}$. Moreover, since $p$ is a vertex we obtain that $\tilde{q}=\tilde{s}$. Therefore $q=i_{\ast}(\tilde{q})=i_{\ast}(\tilde{s})=s$.
Now, suppose that $i_{\ast}(p)$ is a vertex. By Proposition 5.3 from [7] the map $i_{\ast}$ is convex. By Proposition 5.15 [10] every vertex in $(i_{\ast})^{-1}(i_{\ast}(p))$ is a vertex in $\operatorname{sDist}(f)$. Since $i_{\ast}$ is injective we obtain that $(i_{\ast})^{-1}(i_{\ast}(p))=\{p\}$. Therefore $p$ is a vertex.
Part (3): Follows directly from parts (1)-(2) and the fact that a vertex can either be deterministic or contextual. ∎
4 Cycle scenario with arbitrary outcome
In this section, we prove our main result on the classification of the contextual vertices of the polytope of simplicial distributions in cycle scenarios with arbitrary outcomes.
Definition 4.1.
Let $1\leq k\leq d$. A simplicial distribution $p:C^{(n)}\to D(\Delta_{\mathbb{Z}_{d}})$ is called a $k$-order cycle distribution if there exists a finite sequence $\vec{a}=(a_{1}^{(1)},\cdots,a_{n}^{(1)};a_{1}^{(2)},\cdots,a_{n}^{(2)};\cdots;a_{1}^{(k)},\cdots,a_{n}^{(k)})$ of elements in $\mathbb{Z}_{d}$ such that $a_{i}^{(j)}\neq a_{i}^{(s)}$ for every $1\leq i\leq n$, $j\neq s$, and the distribution is defined by
$p_{\sigma_{i}}^{ab}=\left\{\begin{array}[]{ll}\frac{1}{k}&(a,b)=(a_{i}^{(j)},a_{i+1}^{(j)})\;\,\text{for some $1\leq j\leq k$}\\ 0&\text{otherwise,}\end{array}\right.$
for $11\leq i\leq n-$ 1, and
$p_{\sigma_{n}}^{ab}=\left\{\begin{array}[]{ll}\frac{1}{k}&(a,b)=(a_{n}^{(j)},a_{1}^{(j+1)})\;\,\text{for some $1\leq j\leq k-1$}\\ \frac{1}{k}&(a,b)=(a_{n}^{(k)},a_{1}^{(1)})\\ 0&\text{otherwise.}\end{array}\right.$
Note that $1$-order cycle distributions coincide with deterministic distributions.
Example 4.2.
The distribution $p:C^{(2)}\to D(\Delta_{\mathbb{Z}_{4}})$ defined by
$p_{\sigma_{1}}=\begin{pmatrix}0&\frac{1}{3}&0&0\\ 0&0&0&0\\ 0&0&0&\frac{1}{3}\\ 0&0&\frac{1}{3}&0\\ \end{pmatrix}\;\;\text{ and }\;\;p_{\sigma_{2}}=\begin{pmatrix}0&0&0&0\\ 0&0&0&\frac{1}{3}\\ 0&0&\frac{1}{3}&0\\ \frac{1}{3}&0&0&0\\ \end{pmatrix}$
is a $3$-order cycle distribution that comes from the sequence $\vec{a}=(0,1;3,2;2,3)$.
Our key observation is that the $k$-order cycle distributions are contextual vertices that are detected homotopically. For this we rely on the homotopical tools developed in Sections 2.3 and 2.4.
Proposition 4.3.
For $k\geq 2$, every $k$-order cycle distribution is a contextual vertex.
Proof.
We define a simplicial map $\varphi:C^{(n)}\to N\mathbb{Z}_{k}$ by setting
$\varphi_{\sigma_{i}}=\left\{\begin{array}[]{ll}1&i=n\\ 0&\text{otherwise}.\end{array}\right.$
According to Proposition 2.11 the map $\varphi$ is not null-homotopic. Let $p$ be a distribution in $\operatorname{Face}(\varphi)$, i.e., $D(\kappa)(p)=\delta^{\varphi}$. For $11\leq i\leq n-$ 1 we have $D(\kappa)(p)_{\sigma_{i}}=\delta^{0}$, i.e., $\sum_{a=0}^{k-1}p_{\sigma_{i}}^{aa}=1$, which means that $p_{\sigma_{i}}^{ab}=0$ if $b\neq a$. In addition, $D(\kappa)(p)_{\sigma_{n}}=\delta^{1}$, and thus $\sum_{a=0}^{k-1}p_{\sigma_{n}}^{a(a+1)}=1$, that means $p_{\sigma_{n}}^{ab}=0$ if $b\neq a+1$. So
$p^{b}_{d_{0}(\sigma_{i})}=\sum_{a=0}^{k-1}p_{\sigma_{i}}^{ab}=p_{\sigma_{i}}^{bb}\;\;\text{and}\;\;p^{a}_{d_{1}(\sigma_{i})}=\sum_{b=0}^{k-1}p_{\sigma_{i}}^{ab}=p_{\sigma_{i}}^{aa},$
whereas we also have
$p^{b}_{d_{0}(\sigma_{n})}=\sum_{a=0}^{k-1}p_{\sigma_{n}}^{ab}=p_{\sigma_{n}}^{(b-1)b}\;\;\text{and}\;\;p^{a}_{d_{1}(\sigma_{n})}=\sum_{b=0}^{k-1}p_{\sigma_{n}}^{ab}=p_{\sigma_{n}}^{a(a+1)}.$
Therefore
$p_{\sigma_{1}}^{00}=p_{d_{0}(\sigma_{1})}^{0}=p_{d_{1}(\sigma_{2})}^{0}=p_{\sigma_{2}}^{00}=\cdots=p_{\sigma_{n-1}}^{00}=p_{d_{0}(\sigma_{n-1})}^{0}=p_{d_{1}(\sigma_{n})}^{0}=p_{\sigma_{n}}^{01}$
and $p_{\sigma_{n}}^{01}=p_{d_{0}(\sigma_{n})}^{1}=p_{d_{1}(\sigma_{1})}^{1}=p_{\sigma_{1}}^{11}$. Similarly, we have
$p_{\sigma_{1}}^{11}=p_{d_{0}(\sigma_{1})}^{1}=p_{d_{1}(\sigma_{2})}^{1}=p_{\sigma_{2}}^{11}=\cdots=p_{\sigma_{n-1}}^{11}=p_{d_{0}(\sigma_{n-1})}^{1}=p_{d_{1}(\sigma_{n})}^{1}=p_{\sigma_{n}}^{12}$
and $p_{\sigma_{n}}^{12}=p_{d_{0}(\sigma_{n})}^{2}=p_{d_{1}(\sigma_{1})}^{2}=p_{\sigma_{1}}^{22}$. At the end, we obtain that
$p_{\sigma_{1}}^{(k-1)(k-1)}=p_{d_{0}(\sigma_{1})}^{k-1}=p_{d_{1}(\sigma_{2})}^{k-1}=p_{\sigma_{2}}^{(k-1)(k-1)}=\cdots=p_{\sigma_{n-1}}^{(k-1)(k-1)}=p_{d_{0}(\sigma_{n-1})}^{k-1}=p_{d_{1}(\sigma_{n})}^{k-1}=p_{\sigma_{n}}^{(k-1)0}$
and $p_{\sigma_{n}}^{(k-1)0}=p_{d_{0}(\sigma_{n})}^{0}=p_{d_{1}(\sigma_{1})}^{0}=p_{\sigma_{1}}^{00}$. We conclude that $p_{\sigma_{i}}^{aa}=\frac{1}{k}$ for every $11\leq i\leq n-$ 1 and $p_{\sigma_{n}}^{a(a+1)}=\frac{1}{k}$. We proved that the only distribution in $\operatorname{Face}(\varphi)$ is the $k$-cycle distribution that corresponds to the following sequence
$(0,\cdots,0;1,\cdots,1;\cdots;k-1,\cdots,k-1).$
Let us denote this distribution by $q$. By Corollary 2.14 we obtain that $q$ is a contextual vertex. Now, given a $k$-order cycle distribution $p:C^{(n)}\to D(\Delta_{\mathbb{Z}_{d}})$ as in Definition 4.1. We define $T:C^{(n)}\times\Delta_{\mathbb{Z}_{k}}\to C^{(n)}\times\Delta_{\mathbb{Z}_{d}}$ by setting
$T(v_{i},j)=(v_{i},t_{i}(j))$
where $t_{i}(j)=a_{i}^{(j)}$ (see Example 3.2). The simplicial map $T$ is injective since $a_{i}^{(j)}\neq a_{i}^{(r)}$ for every $1\leq i\leq n$, $j\neq r$. Therefore by part (3) of Proposition 3.5 we obtain that the distribution $p=T_{\ast}(q)$ is a contextual vertex. ∎
The converse of this result is also true, leading to the main result of our paper, which characterizes the vertices of cycle scenarios with arbitrary outcomes. For the converse, our main tool is the characterization of vertices using the preorder introduced in Section 2.2.
Theorem 4.4.
A simplicial distribution
$p:C^{(n)}\to D(\Delta_{\mathbb{Z}_{d}})$
is a contextual vertex if and only if it is a $k$-order cycle distribution for some $k\geq 2$.
Proof.
Proposition 4.3 shows one direction. For the other direction, let $p:C^{(n)}\to D(\Delta_{\mathbb{Z}_{d}})$ be a contextual vertex. There is $a_{1},a_{2}\in\mathbb{Z}_{d}$ such that $p_{\sigma_{1}}^{a_{1}a_{2}}\neq 0$, which means that $p_{d_{1}(\sigma_{2})}^{a_{2}}=p_{d_{0}(\sigma_{1})}^{a_{2}}\neq 0$. Therefore there exists $a_{3}\in\mathbb{Z}_{d}$ such that $p_{\sigma_{2}}^{a_{2}a_{3}}\neq 0$. By this process we obtain $a_{1},a_{2},\cdots,a_{n},a_{n+1}$ such that $p_{\sigma_{i}}^{a_{i}a_{i+1}}\neq 0$ for all $1\leq i\leq n$. If $a_{n+1}=a_{1}$ we can define the simplicial map $\varphi:C^{(n)}\to\Delta_{\mathbb{Z}_{d}}$ by setting $\varphi^{\sigma_{i}}=(a_{i},a_{i+1})$ for $1\leq i\leq n$. Then $\varphi\in\operatorname{supp}(p)$, but by part (1) of Proposition 2.7 $p$ is strongly contextual. Therefore $a_{n+1}\neq a_{1}$ and we can continue the process. Then for some $k\geq 2$ we get the following sequence
$a_{1}^{(1)},\cdots,a_{n}^{(1)};a_{1}^{(2)},\cdots,a_{n}^{(2)};\cdots;a_{1}^{(k)},\cdots,a_{n}^{(k)}\in\mathbb{Z}_{d}$
such that
  1. 1.
    $a_{i}^{(j)}\neq a_{i}^{(s)}$ for $1\leq i\leq n$, $j\neq s$,
  2. 2.
    $p_{\sigma_{i}}^{a_{i}^{(j)}a_{i+1}^{(j)}}\neq 0$ for $11\leq i\leq n-$ 1, $1\leq j\leq k$,
  3. 3.
    $p_{\sigma_{n}}^{a_{n}^{(j)}a_{1}^{(j+1)}}\neq 0$ for $11\leq j\leq k-$ 1, and $p_{\sigma_{n}}^{a_{n}^{(k)}a_{1}^{(1)}}\neq 0$.
Let $q$ be the corresponding $k$-order cycle distribution for the sequence above (see Definition 4.1). We conclude that $q\preceq p$. Then by Corollary 2.6 we obtain that $p=q$. ∎
Consider the bundle scenario $f_{m}:E(C^{(n)},\vec{m})\to C^{(n)}$ in Example 3.1. By setting $m=\max\{m_{x}:\,x\in C^{(n)}_{0}\}$ we obtain a bundle morphism
[…]
where $i$ is injective.
Definition 4.5.
We call a simplicial distribution $p$ on the bundle scenario $f_{m}$ a $k$-order cycle distribution if $i_{\ast}(p)$ is a $k$-order cycle distribution in the sense of Definition 4.1.
By Theorem 4.4 and part $(3)$ of Proposition 3.5 we obtain the following result.
Corollary 4.6.
A distribution on the bundle scenario $f_{m}$ is a contextual vertex if and only if it is a $k$-order cycle distribution for some $k\geq 2$.
Corollary 4.6 generalizes Theorem 1 from [8]. In that reference the authors describe the contextual (non-local) extremal distributions (vertices) on the $4$-circle. Our result above extends this characterization to $n$-circle spaces with arbitrary outcomes.
5 Scenarios obtained by gluing
In the theory of simplicial distributions an important strategy is to decompose a measurement space into smaller pieces. In this section we utilize this strategy to provide a characterization which tells us when a simplicial distribution on a union of two measurement spaces is a vertex. Then we apply this characterization to describe vertices of scenarios obtained by gluing cycle scenarios.
We begin by an observation on the relationship between convex decompositions and the preorder on simplicial distributions.
Lemma 5.1.
If $q\in\operatorname{conv}(\operatorname{Vsupp}(p))$, then $q\preceq p$.
Proof.
Assume that $q\in\operatorname{conv}(\operatorname{Vsupp}(p))$. This means that there exists $\alpha_{1},\cdots,\alpha_{n}$ and $q_{1},\cdots,q_{n}\in\operatorname{Vsupp}(p)$, where $\sum_{i=1}^{n}\alpha_{i}=1$, such that $q=\alpha_{1}q_{1}+\cdots+\alpha_{n}q_{n}$. So if $q_{x}^{y}\neq 0$ for some $x\in X_{n}$ and $y\in Y_{n}$, then there exists $1\leq j\leq n$ such that $(q_{j})_{x}^{y}\neq 0$. This implies that $p_{x}^{y}\neq 0$. ∎
Our decomposition result relies on this basic observation.
Proposition 5.2.
Let $X=A\cup B$ and $p:X\to D(Y)$ be a simplicial distribution. There is a bijective correspondence between
  1. 1.
    the set of simplicial distributions $q$ on $X$ satisfying $q\preceq p$, and
  2. 2.
    the set of pairs $(q_{A},q_{B})$ of simplicial distributions $q_{A}\in\operatorname{conv}(\operatorname{Vsupp}(p|_{A}))$ and ${q_{B}\in}\operatorname{conv}(\operatorname{Vsupp}(p|_{B}))$ satisfying $q_{A}|_{A\cap B}=q_{B}|_{A\cap B}$.
Proof.
Let $q:X\to D(Y)$ be such that $q\preceq p$. Then $q|_{A}\preceq p|_{A}$, and because the relation $\preceq$ is transitive we have that $\operatorname{Vsupp}(q|_{A})\subseteq\operatorname{Vsupp}(p|_{A})$. This implies that $q|_{A}\in\operatorname{conv}(\operatorname{Vsupp}(p|_{A}))$. Similarly, we have that $q|_{B}\in\operatorname{conv}(\operatorname{Vsupp}(p|_{B}))$.
Conversely, let $q_{A}$ and $q_{B}$ be such that $q_{A}\in\operatorname{conv}(\operatorname{Vsupp}(p|_{A}))$, $q_{B}\in\operatorname{conv}(\operatorname{Vsupp}(p|_{B}))$, and $q_{A}|_{A\cap B}=q_{B}|_{A\cap B}$. We can construct $q$ on $X$ by gluing $q_{A}$ and $q_{B}$. By Lemma 5.1 we have $q|_{A}\preceq p|_{A}$ and $q|_{B}\preceq p|_{B}$, which implies that $q\preceq p$. ∎
For us this result will be most useful in vertex detection.
Theorem 5.3.
Let $X=A\cup B$. A simplicial distribution $p:X\to D(Y)$ is a vertex if and only if $p$ is the unique simplicial distribution whose restrictions to $A$ and $B$ fall inside $\operatorname{conv}(\operatorname{Vsupp}(p|_{A}))$ and $\operatorname{conv}(\operatorname{Vsupp}(p|_{B}))$, respectively.
Proof.
Follows immediately from Corollary 2.6 and Proposition 5.2. ∎
Next, we apply this technique to identify the vertices of some common measurement scenarios, such as the cycle and Bell scenarios.
5.1 Popescu-Rohrlich boxes
We begin by applying our techniques to a well-known case: The contextual vertices on the scenario $(C^{(4)},\Delta_{\mathbb{Z}_{2}})$ consist of the so-called Popescu-Rohrlich (PR) boxes [14]. Our techniques can be used to prove that PR boxes are vertices. An example of a PR box is given by the distribution
$p|_{\sigma_{1}}=\begin{pmatrix}0&\frac{1}{2}\\ \frac{1}{2}&0\end{pmatrix}\quad\text{and}\quad p|_{\sigma_{i}}=\begin{pmatrix}\frac{1}{2}&0\\ 0&\frac{1}{2}\end{pmatrix}$ (5)
where $i=2,3,4$. Let us consider the decomposition $C^{(4)}=A\cup B$, where
  • $A$ consists of a single edge ($1$-simplex) $\sigma_{1}$, and
  • $B$ consists of the union $\sigma_{2}\cup\sigma_{3}\cup\sigma_{4}$ of three edges; see Figure 2.
Figure 2: Decomposition of the $4$-circle.
Since $A$ is a $1$-simplex and $B$ is a directed path obtained by gluing along 00-simplices it is known (see, e.g., [1, Example 3.11, Corollary 4.6]) that any distribution on $(A,\Delta_{\mathbb{Z}_{d}})$ and $(B,\Delta_{\mathbb{Z}_{d}})$ is noncontextual, thus the vertices are all deterministic. It follows that $\operatorname{Vsupp}(p|_{K}){\cong}\operatorname{supp}(p|_{K})$ for $K=A,B$. So one can see that
$\operatorname{Vsupp}(p|_{A})$ $=$ $\left\{\delta^{\varphi_{1}^{A}},\delta^{\varphi_{2}^{A}}\right\}\quad\text{where}\quad\delta^{\varphi_{1}^{A}}=\begin{pmatrix}0&1\\ 0&0\end{pmatrix},~{}\delta^{\varphi_{2}^{A}}=\begin{pmatrix}0&0\\ 1&0\end{pmatrix}$ (6)
and
$\operatorname{Vsupp}(p|_{B})$ $=$ $\left\{\delta^{\varphi_{1}^{B}},\delta^{\varphi_{2}^{B}}\right\}\quad\text{where}\quad\delta^{\varphi_{1}^{B}}|_{\sigma{i}}=\begin{pmatrix}1&0\\ 0&0\end{pmatrix},~{}\delta^{\varphi_{2}^{B}}|_{\sigma{i}}=\begin{pmatrix}0&0\\ 0&1\end{pmatrix},$ (7)
where $i=2,3,4$.
Let $q_{A}$ and $q_{B}$ be distributions in $\operatorname{conv}\left(\operatorname{Vsupp}(p|_{A})\right)$ and $\operatorname{conv}\left(\operatorname{Vsupp}(p|_{B})\right)$, respectively. Since there are two vertices in the vertex support, these distributions can be written as
$q_{A}=\begin{pmatrix}0&\alpha\\ 1-\alpha&0\end{pmatrix}\quad\quad\text{and}\quad\quad q_{B}|_{\sigma_{i}}=\begin{pmatrix}\beta&0\\ 0&1-\beta\end{pmatrix}$ (8)
for $i=2,3,4$ and $\alpha,\beta\in[0,1]$. Requiring compatibility, i.e.,
${q_{A}|_{A\cap B}=q_{B}|_{A\cap B},}$
gives us that $\alpha=\beta=1-\alpha=1/2$. Since this yields a unique distribution (which is the PR box in Equation (5)), by Theorem 5.3 we have that the PR box is a vertex of $\operatorname{sDist}(C^{(4)},\Delta_{\mathbb{Z}_{2}})$.
5.2 A scenario with trichotomic measurements
We consider a simplicial distribution $p:X\to D(\Delta_{\mathbb{Z}_{3}})$ where $X$ is a $1$-dimensional space depicted in Figure 3. The generating simplices $\sigma_{i}$ $(i=1,\cdots,4)$ have distributions $p|_{\sigma_{i}}=Q_{i}$ where
$Q_{1}=\begin{pmatrix}0&\frac{1}{3}&0\\ \frac{1}{3}&0&0\\ 0&0&\frac{1}{3}\end{pmatrix},~{}~{}Q_{2}=\begin{pmatrix}\frac{1}{3}&0&0\\ 0&\frac{1}{3}&0\\ 0&0&\frac{1}{3}\end{pmatrix},~{}~{}Q_{3}=\begin{pmatrix}\frac{1}{3}&0&0\\ 0&\frac{1}{3}&0\\ 0&0&\frac{1}{3}\end{pmatrix},~{}~{}Q_{4}=\begin{pmatrix}0&0&\frac{1}{3}\\ 0&\frac{1}{3}&0\\ \frac{1}{3}&0&0\end{pmatrix}.$ (9)
We decompose our space into $X=A\cup B$ where $A$ consists of the union $\sigma_{1}\cup\sigma_{2}$ and $B$ consists of $\sigma_{3}\cup\sigma_{4}$. Using Theorem 4.4 we find the vertex supports
$\operatorname{Vsupp}(p|_{A})=\left\{\begin{matrix}\begin{pmatrix}0&0&0\\ 0&0&0\\ 0&0&1\end{pmatrix}_{\sigma_{1}}\\ \begin{pmatrix}0&0&0\\ 0&0&0\\ 0&0&1\end{pmatrix}_{\sigma_{2}}\end{matrix},~{}\begin{matrix}\begin{pmatrix}0&\frac{1}{2}&0\\ \frac{1}{2}&0&0\\ 0&0&0\end{pmatrix}_{\sigma_{1}}\\ \begin{pmatrix}\frac{1}{2}&0&0\\ 0&\frac{1}{2}&0\\ 0&0&0\end{pmatrix}_{\sigma_{2}}\end{matrix}\right\},~{}\operatorname{Vsupp}(p|_{B})=\left\{\begin{matrix}\begin{pmatrix}0&0&0\\ 0&1&0\\ 0&0&0\end{pmatrix}_{\sigma_{3}}\\ \begin{pmatrix}0&0&0\\ 0&1&0\\ 0&0&0\end{pmatrix}_{\sigma_{4}}\end{matrix},~{}\begin{matrix}\begin{pmatrix}\frac{1}{2}&0&0\\ 0&0&0\\ 0&0&\frac{1}{2}\end{pmatrix}_{\sigma_{3}}\\ \begin{pmatrix}0&0&\frac{1}{2}\\ 0&0&0\\ \frac{1}{2}&0&0\end{pmatrix}_{\sigma_{4}}\end{matrix}\right\}.$ (10)
Then arbitrary distributions $q_{A}$ and $q_{B}$ in $\operatorname{conv}\left(\operatorname{Vsupp}(p|_{A})\right)$ and $\operatorname{conv}\left(\operatorname{Vsupp}(p|_{B})\right)$, respectively, are given by
$q_{A}=\begin{matrix}\begin{pmatrix}0&\frac{\alpha_{2}}{2}&0\\ \frac{\alpha_{2}}{2}&0&0\\ 0&0&\alpha_{1}\end{pmatrix}_{\sigma_{1}}\\ \begin{pmatrix}\frac{\alpha_{2}}{2}&0&0\\ 0&\frac{\alpha_{2}}{2}&0\\ 0&0&\alpha_{1}\end{pmatrix}_{\sigma_{2}}\end{matrix}\quad\quad\text{and}\quad\quad q_{B}=\begin{matrix}\begin{pmatrix}\frac{\beta_{2}}{2}&0&0\\ 0&\beta_{1}&0\\ 0&0&\frac{\beta_{2}}{2}\end{pmatrix}_{\sigma_{3}}\\ \begin{pmatrix}0&0&\frac{\beta_{2}}{2}\\ 0&\beta_{1}&0\\ \frac{\beta_{2}}{2}&0&0\end{pmatrix}_{\sigma_{4}}\end{matrix}.$ (11)
The distributions agree on their intersection if and only if
$\frac{1}{2}\alpha_{2}=\frac{1}{2}\beta_{2}\quad\quad\text{and}\quad\quad\frac{1}{2}\alpha_{2}=\beta_{1}.$ (12)
This implies that $\beta_{1}=\beta_{2}/2$. Using normalization $\beta_{1}+\beta_{2}=1$ we obtain $\beta_{1}=1/3$ and $\beta_{2}=2/3$, which additionally yields $\alpha_{1}=1/3$ and $\alpha_{2}=2/3$. This is a unique solution that recovers the distribution $p$ given above. Thus $p$ is a vertex by Theorem 5.3.
Figure 3: The measurement space decomposed into two parts given by $2$-circles.
5.3 $(2,3,3)$ Bell scenario
Next we prove that the distribution on the $(2,3,3)$ Bell scenario, i.e., two parties three trichotomic measurements per party, described in Table III of [15] is a contextual vertex. Consider the simplicial scenario $(X,\Delta_{\mathbb{Z}_{3}})$ where $X$ is a $1$-dimensional space whose underlying graph is $K_{3,3}$; see Figure 4. This corresponds to the bipartite $(2,3,3)$ Bell scenario with three measurements per party and three outcomes per measurement. There are $9$ generating simplices of $X$ that we label $\sigma_{i}$ $(i=1,\cdots,9)$. Introduce now a collection $P,Q,R,S,T$ of distributions on $(\Delta^{1},\Delta_{\mathbb{Z}_{3}})$ given by
$P=\begin{pmatrix}\frac{1}{4}&0&\frac{1}{4}\\ 0&\frac{1}{4}&0\\ \frac{1}{4}&0&0\end{pmatrix},\,Q=\begin{pmatrix}\frac{1}{4}&\frac{1}{4}&0\\ \frac{1}{4}&0&0\\ 0&0&\frac{1}{4}\end{pmatrix},\,R=\begin{pmatrix}\frac{1}{4}&0&\frac{1}{4}\\ \frac{1}{4}&0&0\\ 0&\frac{1}{4}&0\end{pmatrix},\,S=\begin{pmatrix}0&\frac{1}{4}&\frac{1}{4}\\ \frac{1}{4}&0&0\\ \frac{1}{4}&0&0\end{pmatrix},\,T=\begin{pmatrix}\frac{1}{2}&0&0\\ 0&\frac{1}{4}&0\\ 0&0&\frac{1}{4}\end{pmatrix}.$ (13)
Figure 4:
We consider a simplicial distribution $p:X\to\Delta_{\mathbb{Z}_{3}}$ such that $p_{\sigma_{1}}=P$, $p_{\sigma_{2}}=Q$, $p_{\sigma_{3}}=p_{\sigma_{4}}=R$, $p_{\sigma_{5}}=S$, and $p_{\sigma_{i}}=T$ for $i=6,\cdots,9$ (see Figure 4). Notice that the diagonal of $T$ sums to one, thus we can interpret $T$ as a collapsed distribution in the sense of [4, Section $5$]. This means that the distribution $T$ on the edge actually comes from a point via a collapsing map. To see this let $q:\Delta^{0}\to D(\Delta_{\mathbb{Z}_{3}})$ be the distribution on a point where $q=(1/2,1/4,1/4)$. Then we have the diagram
[…]
where $s^{0}$ is the collapsing map which maps an edge to the point. Thus we can collapse the edges in $X$ with distributions $p_{\sigma_{i}}=T$, yielding the quotient space $\bar{X}$ depicted in Figure 5(a). More explicitly, we have the collapse map $\pi:X\to\bar{X}$ which induces a map $\pi^{\ast}:\operatorname{sDist}(\bar{X},\Delta_{\mathbb{Z}_{3}})\to\operatorname{sDist}(X,\Delta_{\mathbb{Z}_{3}})$. In Figure 5(a) we describe the distribution $\bar{p}\in\operatorname{sDist}(\bar{X},\Delta_{\mathbb{Z}_{3}})$ satisfying $\pi^{\ast}(\bar{p})=p$. By part 3 of Theorem 4.4 in [6] we have that $p$ is a vertex of $\operatorname{sDist}(X,\Delta_{\mathbb{Z}_{3}})$ if and only if $\bar{p}$ is a vertex of $\operatorname{sDist}(\bar{X},\Delta_{\mathbb{Z}_{3}})$. Moreover, it is clear that $\bar{p}$ is a vertex if and only if $q:Z\to\Delta_{\mathbb{Z}_{3}}$ is a vertex, where $Z$ and the distribution $q$ are given as in Figure 5(b).
Figure 5: (a) Space obtained from $K_{3,3}$ by collapsing the edges with distribution $T$ to a point. (b) One of the two copies of the edge with distribution $R$ is removed. (c) Measurement space decomposed into two parts.
We are now in a position to apply Theorem 5.3. Consider the decomposition $Z=A\cup B$, where each piece is a $2$-circle $C^{(2)}$ as in Figure 5(c). The corresponding vertex supports are given by
$\operatorname{Vsupp}(q|_{A})$ $=$ $\left\{\begin{matrix}\begin{pmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix}_{\sigma_{1}}\\ \begin{pmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix}_{\sigma_{2}}\end{matrix},~{}\begin{matrix}\begin{pmatrix}\frac{1}{2}&0&0\\ 0&\frac{1}{2}&0\\ 0&0&0\end{pmatrix}_{\sigma_{1}}\\ \begin{pmatrix}0&\frac{1}{2}&0\\ \frac{1}{2}&0&0\\ 0&0&0\end{pmatrix}_{\sigma_{2}}\end{matrix},~{}\begin{matrix}\begin{pmatrix}0&0&\frac{1}{3}\\ 0&\frac{1}{3}&0\\ \frac{1}{3}&0&0\end{pmatrix}_{\sigma_{1}}\\ \begin{pmatrix}0&\frac{1}{3}&0\\ \frac{1}{3}&0&0\\ 0&0&\frac{1}{3}\end{pmatrix}_{\sigma_{2}}\end{matrix},~{}\begin{matrix}\begin{pmatrix}0&0&\frac{1}{2}\\ 0&0&0\\ \frac{1}{2}&0&0\end{pmatrix}_{\sigma_{1}}\\ \begin{pmatrix}\frac{1}{2}&0&0\\ 0&0&0\\ 0&0&\frac{1}{2}\end{pmatrix}_{\sigma_{2}}\end{matrix}\right\}$ (14)
$\operatorname{Vsupp}(q|_{B})$ $=$ $\left\{\begin{matrix}\begin{pmatrix}0&0&0\\ 1&0&0\\ 0&0&0\end{pmatrix}_{\sigma_{4}}\\ \begin{pmatrix}0&1&0\\ 0&0&0\\ 0&0&0\end{pmatrix}_{\sigma_{5}}\end{matrix}~{},\begin{matrix}\begin{pmatrix}0&0&1\\ 0&0&0\\ 0&0&0\end{pmatrix}_{\sigma_{4}}\\ \begin{pmatrix}0&0&0\\ 0&0&0\\ 1&0&0\end{pmatrix}_{\sigma_{5}}\end{matrix}~{},\begin{matrix}\begin{pmatrix}\frac{1}{2}&0&0\\ 0&0&0\\ 0&\frac{1}{2}&0\end{pmatrix}_{\sigma_{4}}\\ \begin{pmatrix}0&0&\frac{1}{2}\\ \frac{1}{2}&0&0\\ 0&0&0\end{pmatrix}_{\sigma_{5}}\end{matrix}~{}\right\}.$ (15)
General distributions $\rho_{A}$ and $\rho_{B}$ on the faces defined by $\operatorname{Vsupp}(q|_{A})$ and $\operatorname{Vsupp}(q|_{B})$, respectively, are given by
$\rho_{A}=\begin{matrix}\begin{pmatrix}\alpha_{1}+\frac{\alpha_{2}}{2}&0&\frac{\alpha_{3}}{3}+\frac{\alpha_{4}}{2}\\ 0&\frac{\alpha_{2}}{2}+\frac{\alpha_{3}}{3}&0\\ \frac{\alpha_{3}}{3}+\frac{\alpha_{4}}{2}&0&0\end{pmatrix}_{\sigma_{1}}\\ \begin{pmatrix}\alpha_{1}+\frac{\alpha_{4}}{2}&\frac{\alpha_{2}}{2}+\frac{\alpha_{3}}{3}&0\\ \frac{\alpha_{2}}{2}+\frac{\alpha_{3}}{3}&0&0\\ 0&0&\frac{\alpha_{3}}{3}+\frac{\alpha_{4}}{2}\end{pmatrix}_{\sigma_{2}}\end{matrix}\quad\text{and}\quad\rho_{B}=\begin{matrix}\begin{pmatrix}\frac{\beta_{3}}{2}&0&\beta_{2}\\ \beta_{1}&0&0\\ 0&\frac{\beta_{3}}{2}&0\end{pmatrix}_{\sigma_{4}}\\ \begin{pmatrix}0&\beta_{1}&\frac{\beta_{3}}{2}\\ \frac{\beta_{3}}{2}&0&0\\ \beta_{2}&0&0\end{pmatrix}_{\sigma_{5}}\end{matrix},$ (16)
where $\alpha_{i},\beta_{j}\in[0,1]$ and $\sum_{i=1}^{4}\alpha_{i}=\sum_{j=1}^{3}\beta_{j}=1$. The two distributions agree on their intersection if and only if the following system of equations holds
$\beta_{1}=\frac{1}{2}\alpha_{2}+\frac{1}{3}\alpha_{3},\quad\quad\beta_{3}=\alpha_{2}+\frac{2}{3}\alpha_{3},\quad\quad\beta_{3}=\frac{2}{3}\alpha_{3}+\alpha_{4},\quad\quad\beta_{2}=\frac{1}{3}\alpha_{3}+\frac{1}{2}\alpha_{4}.$ (17)
Setting the middle two equations equal implies that $\alpha_{2}=\alpha_{4}$ that in turn implies that $\beta_{1}=\beta_{2}=\beta_{3}/2$, from which normalization gives $\beta_{1}=\beta_{2}=1/4$ and $\beta_{3}=1/2$. We substitute these into Equation (17) to obtain that
$\frac{\alpha_{2}}{2}+\frac{\alpha_{3}}{3}=\frac{\alpha_{3}}{3}+\frac{\alpha_{4}}{2}=\frac{\alpha_{3}}{3}+\frac{\alpha_{4}}{2}=1/4.$
Together with the normalization we obtain that $\rho_{A}|_{\sigma_{1}}=P$ and $\rho_{A}|_{\sigma_{2}}=Q$. The uniqueness of the solution implies that $q$ is a vertex by Theorem 5.3, which then implies that $p$ is a vertex.
6 Conclusion
We solve the vertex enumeration problem for the polytope of simplicial distributions on cycle scenarios with arbitrary outcomes. The simplicial distribution formulation aligns with the usual non-signaling formulation. Our approach uniquely benefits from novel homotopical methods, extending to scenarios formed by gluing cycle scenarios together. This gluing process is a crucial aspect of the theory of simplicial distributions when studying contextual distributions.
Decomposing a measurement space into smaller components is a notable feature of our work that warrants further investigation. As illustrated in Section 5, if the vertex enumeration problem (see, e.g., [16]) is solved for measurement spaces $A$ and $B$ in a given outcome space, this solution can be leveraged to detect vertices in the composite space $X=A\cup B$. Consequently, our results suggest a novel algorithm for enumerating the vertices of polytopes derived from simplicial distributions. Systematizing this procedure to identify additional classes of contextual vertices in other scenarios is a direction for future research.